# Each Rank Could be an Expert: Single-Ranked Mixture of Experts LoRA for Multi-task Learning

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) is widely used for adapting large language models (LLMs) to specific domains due to its efficiency and modularity. Meanwhile, vanilla LoRA struggles with task conflicts in multi-task scenarios. Recent works adopt Mixture of Experts (MoE) by treating each LoRA module as an expert, thereby mitigating task interference through multiple specialized LoRA modules. While effective, these methods often isolate knowledge within individual tasks, failing to fully exploit the shared knowledge across related tasks. In this paper, we establish a connection between single LoRA and multi-LoRA MoE, integrating them into a unified framework. We demonstrate that the dynamic routing of multiple LoRAs is functionally equivalent to rank partitioning and block-level activation within a single LoRA. We further empirically demonstrate that finer-grained LoRA partitioning, within the same total and activated parameter constraints, leads to better performance gains across heterogeneous tasks. Building on these findings, we propose Single-ranked Mixture of Experts LoRA (SMoRA), which embeds MoE into LoRA by treating each rank as an independent expert. With a dynamic rank-wise activation mechanism, SMoRA promotes finer-grained knowledge sharing while mitigating task conflicts. Experiments demonstrate that SMoRA activates fewer parameters yet achieves better performance in multi-task scenarios.
Machine Learning, ICML
1 Introduction
With the remarkable success of LLMs across various domains (Dubey et al., 2024; Liu et al., 2023b), there has been growing interest in fine-tuning them for diverse, task-specific applications (Zhang et al., 2023b). However, full-parameter fine-tuning is often computationally prohibitive, prompting the adoption of Parameter-Efficient Fine-Tuning (PEFT) approaches (Han et al., 2024). Among these, LoRA has gained prominence for its efficiency and modular design (Han et al., 2024; Zhao et al., 2024b).
Despite its advantages, LoRA encounters a critical limitation in multi-task settings: it uniformly activates all parameters during fine-tuning, without decoupling them to adapt to the specific needs of heterogeneous downstream tasks, as shown in Fig.1a). This lack of task-specific adaptation often leads to performance degradation in multi-task scenarios (Tian et al., 2024; Liu et al., 2023a; Zhao et al., 2024c; Feng et al., 2024).
To address this challenge, recent studies have incorporated the MoE framework into LoRA training (Mao et al., 2025; Dou et al., 2023; Liu et al., 2023a; Zadouri et al., 2023). As illustrated in Fig.1b), these approaches construct multiple LoRAs, treating each module as an independent expert. A router is then employed to dynamically activate specific modules, enhancing task adaptation and computational efficiency (Zhu et al., 2023; Li et al., 2024; Feng et al., 2024). However, these methods partition the parameter space into several fixed submatrices, where each block is activated or deactivated as a whole. This block-wise activation is inherently coarse-grained, failing to capture subtle distinctions or shared knowledge across tasks, leading to inefficient utilization of learned knowledge. Furthermore, the rigid division of modules restricts the dynamic combination and decomposition of knowledge, ultimately limiting the model’s expressive power and adaptability in multi-task learning scenarios.
In this paper, we first provide a unifying perspective, demonstrating that multi-LoRA MoE can be represented as a single LoRA, where the rank is partitioned into blocks and each block is activated independently, as shown in Fig. 2 (see §2.2 for details). Through extensive experiments, we further investigate the impact of block granularity on downstream performance under fixed total and activated parameter budgets. Interestingly, we observe that finer-grained partitioning of the rank space leads to significantly better performance, as illustrated in Fig. 3 (see §2.3 for details). Based on these findings, we propose the following key insight:
@@ 1 Introduction
Building on this motivation, we propose Single-ranked Mixture of Experts LoRA (SMoRA), which treats each rank of LoRA as an expert. By incorporating an MoE structure within a single LoRA, SMoRA enables dynamic rank-wise sparse activation, as shown in Fig.1c). This allows the model to dynamically and flexibly combine ranks, addressing the rigid knowledge isolation in MoE-based LoRA and enabling fine-grained task adaptation and knowledge sharing. However, the dynamic activation mechanism in SMoRA necessitates careful management of rank allocation and computational efficiency to fully harness its potential. To address these challenges, we adopt a loss-free load-balancing strategy (Wang et al., 2024), incorporating an expert bias term to ensure balanced rank utilization without modifying the base model. Additionally, we develop a custom CUDA kernel indexed_matmul using TVM (Chen et al., 2018), which performs dynamic matrix multiplications based on router-selected indices, significantly improving computational efficiency and reducing memory overhead.
@@
  • Customized CUDA Kernel: We implemented a custom CUDA kernel, indexed_matmul, to enable sparse activation and computation within LoRA. By leveraging this kernel, we significantly reduce both computational and GPU memory overhead, leading to a more efficient process.
@@ 2.1 Low-Rank Adaptation & LoRA MoE
Extending this idea, LoRA MoE (Dou et al., 2023; Mao et al., 2025; Liu et al., 2023a; Shen et al., 2023; Zhu et al., 2023) incorporates multiple LoRA modules (referred to as ”experts”) and employs an input-dependent router to select or combine these experts dynamically. Formally, the forward pass of LoRA MoE becomes:
@@
Figure 2: Muli-LoRA MoE Equals to Single-LoRA with Block-wise Activation.
@@ 2.2 The Equivalence Between Multi-LoRA MoE and Single-Rank Blockwise Activation
Since the gating vector ${\bm{g}}({\bm{x}})$ in Eq.3 determines the activation pattern of the experts, the term $\tilde{{\bm{B}}}{\bm{G}}({\bm{x}})\tilde{{\bm{A}}}{\bm{x}}$ represents a weighted combination of the original low-rank updates ${\bm{A}}_{i}{\bm{B}}_{i}{\bm{x}}$, with the activation pattern controlled by the gating function ${\bm{g}}({\bm{x}})$. Therefore, the MoE architecture is equivalent to a single LoRA module with a block-wise activation pattern, where each block corresponds to an expert ${\bm{B}}_{i}{\bm{A}}_{i}$, and the gating function ${\bm{g}}({\bm{x}})$ controls which blocks are activated.
Figure 3: Effect of Expert Division Granularity on Average Performance. The x-axis shows different configurations of expert rank, expert count, and activated expert count, with the total number of parameters and activated parameters held constant.
@@ 2.3 Motivation
As shown in Property 2.1, constructing multiple LoRA experts in multi-LoRA can be unified into a single LoRA framework with block-wise activation, essentially corresponding to segmenting the rank parameter within LoRA into smaller blocks. This equivalence motivates us to investigate how varying granularity in the modular decomposition of LoRA can lead to better performance on downstream heterogeneous tasks.
We thus conducted experiments on the FLAN-v2 dataset (Longpre et al., 2023), initializing LoRA with a rank of 64. To evaluate the impact of partitioning granularity, we applied different levels of partitioning while keeping the number of activated ranks constant at 16. From the results shown in Fig.3, we can observe that, with the total parameter count and activated parameters remaining constant, finer granularity in expert partitioning leads to better performance on downstream heterogeneous data. This is because finer-grained expert partitioning enables more precise control over knowledge division and sharing through the router, while also allowing for a greater variety of expert combinations during inference, leading to more flexible parameter allocation and knowledge sharing. This motivates us to embed the MoE structure within LoRA, treating each rank as an expert, and using rank-wise activation to naturally achieve the finest-grained partitioning, thereby balancing task-specific parameters and knowledge sharing.
@@ 3 Methodology
In this section, we introduce the proposed SMoRA which embeds an MoE into the LoRA architecture to dynamically activate ranks within LoRA for achieving a balance between knowledge sharing and preventing task conflict in multi-task learning.
@@ 3.1 Architecture of SMoRA
The overall architecture of SMoRA is illustrated in Fig.4. Building upon pre-trained weights ${\bm{W}}_{0}\in\mathbb{R}^{d\times d}$, SMoRA introduces additional trainable low-rank matrices ${\bm{B}}\in\mathbb{R}^{d\times r}$ and ${\bm{A}}\in\mathbb{R}^{r\times d}$, as in vanilla LoRA, where $r$ denotes the rank. Additionally, SMoRA extends LoRA by treating each rank as an individual ”expert” and incorporating a sparse MoE mechanism. A gating function ${\bm{g}}({\bm{x}})\in\mathbb{R}^{r}$ determines the importance of each rank for a given input ${\bm{x}}$, enabling dynamic sparse activation of the most relevant ranks via top-k routing. The gating matrix ${\bm{G}}({\bm{x}})=\mathrm{diag}({\bm{g}}_{1}({\bm{x}}),{\bm{g}}_{2}({\bm{x}}),\ldots,{\bm{g}}_{r}({\bm{x}}))$ is diagonal, where ${\bm{g}}_{i}({\bm{x}})$ represents the gating score for the $i$-th rank expert, making the model both computationally efficient and expressive. The forward pass of SMoRA is formulated as:
@@
${\bm{g}}({\bm{x}})=\mathrm{Softmax}(\mathrm{TopK}({\bm{x}}{\bm{W}}_{g})),$ (5)
where ${\bm{x}}\in\mathbb{R}^{d}$ is the input, ${\bm{W}}_{g}\in\mathbb{R}^{d\times r}$ is a learnable projection matrix, and $\mathrm{TopK}(\cdot)$ retains only the top-k largest values, setting others to $\infty$- ∞ before applying the Softmax. The sparsity level is controlled by the hyperparameter $k$, which defines the number of active ranks for each token.
@@ 3.2 Load Balance with Loss-Free Balancing
${\bm{g}}({\bm{x}})=\mathrm{Softmax}(\mathrm{TopK}({\bm{x}}{\bm{W}}_{g}+{\bm{b}})),$ (6)
@@
This training-free load-balancing approach aligns well with the philosophy of PEFT methods, which emphasizes modularity by avoiding modifications to the base model’s code. By sidestepping the need for model-specific balancing losses, our method preserves the plug-and-play nature of adapters, enabling them to be directly loaded as parameters into the linear layers of the model without altering the underlying architecture or training process of the base model.
@@
To enhance the efficiency of sparse matrix computations in SMoRA, we developed a custom CUDA kernel, indexed_matmul, using TVM (Chen et al., 2018). This kernel utilizes the top-K indices from the gating function to selectively extract and process relevant rows and columns from the LoRA matrices, thereby substantially reducing computational overhead and memory consumption. By performing only the essential operations and bypassing full matrix multiplication, indexed_matmul achieves superior speed and GPU memory efficiency.
TVM enhances these computations with GPU-specific optimizations, including optimized memory access patterns and effective parallelization. Unlike prior MoE implementations that rely on time-intensive for-loop iterations to process each expert, our approach eliminates these loops by directly targeting non-zero indices. This results in significantly more efficient execution on CUDA-enabled devices, enabling SMoRA to support larger models with higher ranks while mitigating time complexity and memory bottlenecks. Detailed implementation and additional discussion on TVM optimizations can be found in Appendix D.
Table 1: Performance Across Task Clusters in FLAN-v2. The best performance for each task is highlighted in bold, and the second-best performance is underlined. The performance of our proposed SMoRA is highlighted in gray. The full results of each task are shown in Appendix E
@@
Method GSM8K MMLU Law Medical HumanEval
@@
LoRA-64 34.80 21.37 19.35 29.00 18.90 29.88
MoE-Top1 26.46 21.04 23.83 33.00 15.85 30.49
MoE-Top2 31.39 21.52 27.93 28.00 15.24 25.00
MoE-Soft 28.89 18.10 17.60 27.00 15.85 30.97
SMEAR 27.90 17.73 18.33 29.00 16.46 26.83
HydraLoRA 27.14 18.54 21.58 28.00 15.85 28.05
MoSLoRA 33.66 19.56 17.23 27.00 18.29 30.49
SMoRA 34.95 30.61 34.65 41.00 18.90 35.37
@@ 4 Experiments
To validate the efficacy of current PEFT methods in multi-task learning scenarios, we leverage Llama-2-{7b,13b} (Touvron et al., 2023) as the base models and train a range of LoRAs for a spectrum of tasks. We select a portion of the Flan-v2 datasets which contains 48 tasks covering Natural Language Understanding (NLU) and Natural Language Generation (NLG) that can be grouped into 10 distinct task clusters. To further evaluate the effectiveness of our proposed method in domain-specific tasks, we constructed a multi-domain benchmark and tested it using Llama-2-7b. This benchmark spans general language, medical, legal, mathematical, and code generation tasks, following the approach outlined in (Tian et al., 2024). The evaluation set includes: (1) MMLU (Chen et al., 2021) for general ability, (2) Law tasks from MMLU, (3) Medical tasks from MMLU, (4) the GSM8K (Cobbe et al., 2021) evaluation set for mathematics, and (5) HumanEval (Chen et al., 2021) for code generation. Details of the tasks can be found in Appendix A. Additionally, we evaluate the performance of the Llama3-Instruct model (Dubey et al., 2024) after training on code-related datasets, specifically assessing its performance on multiple data analysis tasks using Python libraries. The results of these evaluations are provided in Appendix B.
@@
We compare the proposed method against several state-of-the-art PEFT methods, including (1) LoRA with varying numbers of ranks; (2) LoRA MoE with top-1 and top-2 routing, where each expert is a LoRA with rank 8; (3) LoRA MoE with soft routing, which computes a weighted average of the experts’ outputs based on the gating scores; (4) SMEAR, which performs LoRA fusion based on the gating scores; (5) HydraLoRA, which leverages an asymmetric structure consisting of one ${\bm{A}}$ matrix and four ${\bm{B}}$ matrices, each with rank 8; (6) MoSLoRA, which incorporates a mixture matrix between the LoRA matrices ${\bm{A}}$ and ${\bm{B}}$. Detailed implementations of each baseline method can be found in Appendix C.
@@
The total rank of all LoRA MoE baselines is set to 64. SMoRA follows this setting with a total rank of 64, activating 8 rank parameters with an additional router. The update rate $u$ of the expert routing bias in Eq.6 is set to $1\times 10^{-5}$ for all SMoRA experiments. Additionally, the initialization of SMoRA’s ${\bm{A}}$ and ${\bm{B}}$ matrices is identical to that of vanilla LoRA, ensuring consistency in the base parameter setup.
@@
The main results for Flan-v2 are presented in Tab.1. Based on these results, we make the following observations: (1) The performance of Vanilla LoRA improves gradually as the rank increases. However, due to task interference, it performs worse than the MoE and SMoRA methods when compared under the same number of activated parameters. (2) Due to the partitioning of parameters, MoE methods perform better than LoRA counterparts with the same number of activated parameters. However, as they fail to fully utilize shared knowledge across different tasks, their performance remains suboptimal. (3) By employing rank-wise activation and fine-grained parameter partitioning, SMoRA effectively decouples parameters, striking a balance between task interference and knowledge sharing. This allows SMoRA to achieve the best performance. (4) Notably, SMoRA, which activates only 8 out of 64 ranks in LoRA, outperforms the full Vanilla LoRA model (with all 64 ranks activated), achieving a performance improvement of 1.73% in the Llama2-7b experiment and 1.33% in Llama2-13b. Furthermore, compared to a LoRA model with 8 activated ranks, SMoRA demonstrates an 11.16% improvement in Llama2-7b and 5.34% in Llama2-13b. Additionally, SMoRA outperforms LoRA MoE, which treats every 8 ranks as a separate expert and activates the top-1 expert, achieving a 6.13% improvement in Llama2-7b and 3.74% in Llama2-13b. These results emphasize SMoRA’s superior efficiency and performance, even with a smaller subset of activated ranks. We also compared the activated parameter count across different methods in Appendix H.
Figure 5: Visualization of Routing Distribution. Tasks from the same domain are grouped in square brackets.
Figure 6: The Impact of Load Balancing Strategy on the Expert Load Radio.
@@
To better understand the routing distribution of SMoRA, we visualize it by selecting a model layer and plotting the routing patterns for a portion of the evaluation set, categorized by each task. From Fig. 5a), we observe that rank-wise routing effectively decouples different parameters, with each task exhibiting a distinct expert distribution. This demonstrates that different tasks can activate unique rank parameters, enabling parameter decoupling and reducing task interference. Moreover, similar tasks tend to share comparable routing distributions. As illustrated in Fig. 5b), we plot heatmaps of the routing distributions for various tasks. The heatmaps reveal that tasks within the same task cluster exhibit similar routing patterns, suggesting that SMoRA facilitates improved sharing of related knowledge. This results in fine-grained knowledge sharing and parameter decoupling.

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) is widely used for adapting large language models (LLMs) to specific domains due to its efficiency and modularity. However, vanilla LoRA struggles with task conflicts in multi-task scenarios. Recent works use a Mixture of Experts (MoE) approach, treating each LoRA module as a specialized expert to reduce interference. However, these MoE methods often isolate knowledge within individual tasks, failing to leverage shared knowledge across related tasks. In this paper, we present a unified framework for single LoRA and multi-LoRA MoE. Within this framework, we propose Single-ranked Mixture of Experts LoRA (SMoRA), which embeds MoE into a singular LoRA by treating each rank as an independent expert. With a dynamic rank-wise sparse activation mechanism, SMoRA achieves an automatic rank dispatching that promotes finer-grained knowledge sharing while mitigating task conflicts. Experiments demonstrate that SMoRA activates fewer parameters yet achieves better performance in multi-task scenarios.
Each Rank Could be an Expert: Single-Ranked Mixture of Experts LoRA for Multi-task Learning
Ziyu Zhao12thanks:   Equal Contribution., Yixiao Zhou1211footnotemark: 1, Zhi Zhang3, Didi Zhu1, Tao Shen1, Zexi Li1, Jinluan Yang1, Xuwu Wang3, Jing Su3, Kun Kuang15, Zhongyu Wei4, Fei Wu1 Yu Cheng5 1Zhejiang University, 2Shanghai Innovation Institute, 3ByteDance Inc., 4Fudan University, 5The Chinese University of Hong Kong
With the remarkable success of LLMs across various domains Dubey et al. (2024); Liu et al. (2023b), there has been growing interest in fine-tuning them for diverse, task-specific applications Zhang et al. (2023b). However, full-parameter fine-tuning is often computationally prohibitive, prompting the adoption of Parameter-Efficient Fine-Tuning (PEFT) approaches Han et al. (2024). Among these, LoRA has gained prominence for its efficiency and modular design Han et al. (2024); Zhao et al. (2024b). Despite its advantages, LoRA encounters a critical limitation in multi-task settings: it uniformly activates all parameters during fine-tuning, without decoupling them to adapt to the specific needs of heterogeneous downstream tasks, as shown in Fig.1a). This lack of task-specific adaptation often leads to performance degradation in multi-task scenarios Tian et al. (2024); Liu et al. (2023a); Zhao et al. (2024c); Feng et al. (2024).
@@
To address this challenge, recent studies have incorporated the MoE framework into LoRA training Mao et al. (2025); Dou et al. (2023); Liu et al. (2023a); Zadouri et al. (2023). As illustrated in Fig.1b), these approaches construct multiple LoRAs, treating each module as an independent expert. A router is then employed to dynamically activate specific modules, enhancing task adaptation and computational efficiency Zhu et al. (2023); Li et al. (2024); Feng et al. (2024). However, these methods partition the parameter space into several fixed submatrices, where each block is activated or deactivated as a whole. This block-wise activation is inherently coarse-grained, which fails to capture subtle distinctions or shared knowledge across tasks, leading to inefficient utilization of learned knowledge. Furthermore, the rigid division of modules restricts the dynamic combination and decomposition of knowledge, ultimately limiting the model’s expressive power and adaptability in multi-task learning scenarios.
In this paper, we first provide a unifying perspective, demonstrating that multi-LoRA MoE can be represented as a single LoRA, where the rank is partitioned into blocks and each block is activated independently (see §1.2 for details). Through extensive experiments, we further investigate the impact of block granularity on downstream performance under fixed total and activated parameter budgets. Interestingly, we observe that finer-grained partitioning of the rank space yields significantly better performance and allows the resulting learned experts to represent a larger and more diverse parameter space, as illustrated in Fig. 2 (see §1.3 for details). Based on these findings, we propose the following key insight:
@@
Building on this motivation, we propose Single-ranked Mixture of Experts LoRA (SMoRA), which treats each rank of LoRA as an expert. By incorporating an MoE structure within a single LoRA, SMoRA enables dynamic rank-wise sparse activation, as shown in Fig.1c). This allows the model to dynamically and flexibly combine ranks, addressing the rigid knowledge isolation in MoE-based LoRA and enabling fine-grained task adaptation and knowledge sharing. However, the dynamic activation mechanism in SMoRA necessitates careful management of rank allocation and computational efficiency to fully harness its potential. To address these challenges, we adopt a loss-free load-balancing strategy Wang et al. (2024), incorporating an extra expert bias term to ensure balanced rank utilization without modifying the base model. Additionally, we develop a custom CUDA kernel indexed_matmul using TVM Chen et al. (2018), which performs dynamic matrix multiplications based on router-selected indices, preventing additional computational burden.
@@
  • Unified LoRA Framework and Analysis of MoE Granularity Impact: We establish a unified framework that connects multi-LoRA MoE with standard single-LoRA. Building on this, our in-depth analysis reveals how expert granularity within LoRA-MoE architectures critically impacts model performance and influences the learned parameters’ representational capability.
@@
Figure 2: Varying Expert Granularity. Effects of altering expert granularity while keeping total rank capacity (Expert Rank × Expert Count = 64) and activated rank (Expert Rank × Activate Count = 16) constant. a) Average performance demonstrates an increase as expert granularity becomes finer. b) Rank-wise parameter similarity shifts from block-diagonal structures (coarse-grained) to strong main-diagonal patterns (fine-grained), implying more distinct fine-grained experts capable of representing a larger, more diverse parameter space.
@@ 1.1 Low-Rank Adaptation & LoRA MoE
Extending this idea, LoRA MoE Dou et al. (2023); Mao et al. (2025); Liu et al. (2023a); Shen et al. (2023); Zhu et al. (2023) incorporates multiple LoRA modules (referred to as "experts") and employs an input-dependent router to select or combine these experts dynamically. Formally, the forward pass of LoRA MoE becomes:
@@ 1.2 The Equivalence Between Multi-LoRA MoE and Single-Rank Blockwise Activation
The gating vector in Eq.3 controls expert activation through its diagonal matrix form. This enables the term to function as a block-activated LoRA module, where each expert corresponds to a submatrix block, and its activation is governed by the gating mechanism. In essence, the MoE architecture functions as a single LoRA module with sparse, input-dependent block-wise activation.
@@ 1.3 Motivation
Property 1.1 establishes that constructing multiple LoRA experts is equivalent to a single LoRA framework with block-wise activation, where each expert corresponds to a partition of the rank parameter. This equivalence naturally leads us to question how the granularity of this rank partitioning impacts performance, especially on diverse downstream tasks.
To investigate this, we conducted experiments on the heterogeneous FLAN-v2 dataset Longpre et al. (2023). We initialized LoRA with a total rank of 64 and compared different partitioning schemes while keeping the number of activated ranks constant at 16. As shown in Fig.2 a), our results demonstrate a clear trend: finer partitioning granularity leads to improved performance, even when the total parameter count and the number of activated parameters remain constant.
To understand the underlying structural changes driving this performance gain, in Fig.2 b), we visualized the rank-wise parameter similarity using $[A,B^{T}][A,B^{T}]^{T}$ heatmaps of a layer in LLM. These visualizations reveal distinct structural patterns. Coarse-grained partitioning (e.g., 16 ranks per expert in a 16x4 configuration) results in prominent block-diagonal structures, indicating that the learned parameters within each large expert block are highly similar and thus more concentrated. This internal homogeneity, while ensuring separation between expert blocks, limits the diversity of knowledge each coarse expert can capture, thereby impeding the model’s flexibility in learning varied information. Instead, with fine-grained partitioning (such as each rank-1 component acting as an expert in a 1x64 setup), a prominent main diagonal emerges. This, coupled with increasingly diffuse and complex off-diagonal correlations, points to a substantial reduction in the learned correlation between parameters of different fine-grained experts. These experts thus maintain their distinctiveness rather than forming larger correlated groups, a characteristic that enables the model to capture a wider and more flexible array of knowledge by leveraging more specialized, less redundant expert contributions. Furthermore, this fine-grained structure enables more precise control over knowledge selection and sharing via the routing mechanism and allows for a greater variety of expert combinations during inference. Therefore, these findings motivate our approach: we embed the MoE structure directly within LoRA. Since the parameters of each rank in LoRA represent its smallest constituent unit, we naturally treat each rank as an individual expert. This rank-wise activation achieves the finest possible partitioning granularity, offering a principled way to balance task specialization and knowledge sharing.
@@ 2.1 Architecture of SMoRA
The overall architecture of SMoRA is illustrated in Fig.3. Building upon pre-trained weights ${\bm{W}}_{0}\in\mathbb{R}^{d\times d}$, SMoRA introduces additional trainable low-rank matrices ${\bm{B}}\in\mathbb{R}^{d\times r}$ and ${\bm{A}}\in\mathbb{R}^{r\times d}$, as in vanilla LoRA, where $r$ denotes the rank. Additionally, SMoRA extends LoRA by treating each rank as an individual “expert" and incorporating a sparse MoE mechanism. A gating function ${\bm{g}}({\bm{x}})\in\mathbb{R}^{r}$ determines the importance of each rank for a given input ${\bm{x}}$, enabling dynamic sparse activation of the most relevant ranks via top-k routing. The gating matrix ${\bm{G}}({\bm{x}})=\mathrm{diag}({\bm{g}}_{1}({\bm{x}}),{\bm{g}}_{2}({\bm{x}}),\ldots,{\bm{g}}_{r}({\bm{x}}))$ is diagonal, where ${\bm{g}}_{i}({\bm{x}})$ represents the gating score for the $i$-th rank expert. The forward pass of SMoRA is formulated as:
@@
${\bm{g}}({\bm{x}})=\mathrm{Softmax}(\mathrm{TopK}({\bm{W}}_{g}{\bm{x}})),$ (5)
where ${\bm{x}}\in\mathbb{R}^{d}$ is the input, ${\bm{W}}_{g}\in\mathbb{R}^{r\times d}$ is a learnable projection matrix, and $\mathrm{TopK}(\cdot)$ retains only the top-k largest values, setting others to $\infty$- ∞ before applying the Softmax. The sparsity level is controlled by the hyperparameter $k$, which defines the number of active ranks for each token.
@@ 2.2 Load Balance with Loss-Free Balancing
${\bm{g}}({\bm{x}})=\mathrm{Softmax}(\mathrm{TopK}({\bm{W}}_{g}{\bm{x}}+{\bm{b}})),$ (6)
@@
This training-free load-balancing approach aligns well with the philosophy of PEFT methods, which emphasizes modularity by avoiding modifications to the base model. By sidestepping the need for model-specific balancing losses, our method preserves the plug-and-play nature of adapters, enabling them to be directly loaded as parameters into the linear layers of the model without altering the underlying architecture or training process of the base model.
@@
To improve the efficiency of sparse matrix operations in SMoRA, we developed indexed_matmul, a custom CUDA kernel leveraging TVM Chen et al. (2018). This kernel utilizes top-K indices from the router’s gating function to selectively compute only with the active rows and columns of the LoRA matrices. This targeted computation method primarily avoids two common sources of inefficiency: first, the substantial computational costs inherent in dense matrix operations, by processing only essential data; and second, the latency from iterative processing, as it bypasses the explicit for-loops typically employed in many MoE frameworks to handle individual experts. For a detailed implementation and further discussion of TVM optimizations, please see Appendix F.
Table 1: Performance Across Task Clusters in FLAN-v2. The best performance for each task is highlighted in bold, and the second-best performance is underlined. The performance of our proposed SMoRA is highlighted in gray. The full results of each task are shown in Appendix G
@@
Method GSM8K MMLU Law Medical HumanEval Average
@@
LoRA-64 34.80 41.45 41.79 44.61 18.90 29.88 35.24
MoE-Top1 26.46 38.53 38.41 34.35 15.85 30.49 30.68
MoE-Top2 31.39 39.92 39.65 39.71 15.24 25.00 31.82
MoE-Soft 28.89 41.52 43.98 41.57 15.85 30.97 33.79
SMEAR 27.90 40.68 45.35 38.31 16.46 26.83 32.59
HydraLoRA 27.14 41.54 42.38 39.50 15.85 28.05 32.41
MoSLoRA 33.66 40.65 43.32 41.38 18.29 30.49 34.63
SMoRA 34.95 41.70 45.20 43.81 18.90 35.37 36.66
@@ 3 Experiments
To validate the efficacy of current PEFT methods in multi-task learning scenarios, we leverage Llama-2-{7b,13b} Touvron et al. (2023) as the base models and train a range of LoRAs for a spectrum of tasks. We select a portion of the Flan-v2 datasets, which contains 48 tasks covering Natural Language Understanding (NLU) and Natural Language Generation (NLG), that can be grouped into 10 distinct task clusters. To further evaluate the effectiveness of our proposed method in domain-specific tasks, we constructed a multi-domain benchmark and tested it using Llama-2-7b. This benchmark spans general language, medical, legal, mathematical, and code generation tasks, following the approach outlined in Tian et al. (2024). The evaluation set includes: (1) MMLU Chen et al. (2021) for general ability, (2) Law tasks from MMLU, (3) Medical tasks from MMLU, (4) the GSM8K Cobbe et al. (2021) evaluation set for mathematics, and (5) HumanEval Chen et al. (2021) for code generation. Details of the tasks can be found in Appendix C. Additionally, we evaluate the performance of the Llama3-Instruct model Dubey et al. (2024) after training on code-related datasets, specifically assessing its performance on multiple data analysis tasks using Python libraries. The results of these evaluations are provided in Appendix D.
@@
We compare the proposed method against several state-of-the-art PEFT methods, including (1) LoRA with varying numbers of ranks; (2) LoRA MoE with top-1 and top-2 routing, where each expert is a LoRA with rank 8; (3) LoRA MoE with soft routing, which computes a weighted average of the experts’ outputs based on the gating scores; (4) SMEAR, which performs LoRA fusion based on the gating scores; (5) HydraLoRA, which leverages an asymmetric structure consisting of one ${\bm{A}}$ matrix and four ${\bm{B}}$ matrices, each with rank 8; (6) MoSLoRA, which incorporates a mixture matrix between the LoRA matrices ${\bm{A}}$ and ${\bm{B}}$. Detailed implementations of each baseline method can be found in Appendix E.
@@
The total rank of all LoRA MoE baselines is set to 64. SMoRA follows this setting with a total rank of 64, activating 8 rank parameters with an additional router. The update rate $u$ of the expert routing bias in Eq.6 is set to $1\times 10^{-5}$ for all SMoRA experiments. Additionally, the initialization of SMoRA’s ${\bm{A}}$ and ${\bm{B}}$ matrices is identical to that of vanilla LoRA, ensuring consistency in the base parameter setup. An ablation study on the effect of total activation rank in downstream tasks is provided in Appendix A.
Figure 4: Visualization of Routing Distribution. Tasks from the same domain are grouped in square brackets. The X-axis in Figure a) denotes the expert index.
Figure 5: The Impact of Load Balancing Strategy on Expert Load Distribution. The left figure compares expert load distribution with and without a load balancing strategy, while the right shows the resulting performance differences.
@@
The main results for Flan-v2 are presented in Tab.1. Based on these results, we make the following observations: (1) The performance of Vanilla LoRA improves gradually as the rank increases. However, due to task interference, it performs worse than the MoE and SMoRA methods when compared under the same number of activated parameters. (2) Due to the partitioning of parameters, MoE methods perform better than LoRA counterparts with the same number of activated parameters. However, as they fail to fully utilize shared knowledge across different tasks, their performance remains suboptimal. (3) By employing rank-wise activation and fine-grained parameter partitioning, SMoRA effectively decouples parameters, striking a balance between task interference and knowledge sharing. This allows SMoRA to achieve the best performance. (4) Notably, SMoRA, which activates only 8 out of 64 ranks in LoRA, outperforms the full Vanilla LoRA model (with all 64 ranks activated), achieving a performance improvement of 1.73% in the Llama2-7b experiment and 1.33% in Llama2-13b. Furthermore, compared to a LoRA model with 8 activated ranks, SMoRA demonstrates an 11.16% improvement in Llama2-7b and 5.34% in Llama2-13b. Additionally, SMoRA outperforms LoRA MoE, which treats every 8 ranks as a separate expert and activates the top-1 expert, achieving a 6.13% improvement in Llama2-7b and 3.74% in Llama2-13b. These results emphasize SMoRA’s superior effectiveness, even with a smaller subset of activated ranks. We also compared the activated parameter count across different methods in Appendix J.
@@
To better understand the routing distribution of SMoRA, we visualize it by selecting a model layer and plotting the routing patterns for a portion of the evaluation set, categorized by each task. From Fig.4a), we observe that rank-wise routing effectively decouples parameters, with each task displaying a distinct expert distribution. This indicates that different tasks can activate unique rank parameters, thereby minimizing task interference. Furthermore, tasks with similarities exhibit comparable routing distributions. Fig.4b) presents heatmaps of the routing distributions across various tasks, revealing that tasks within the same cluster share similar routing patterns. This observation suggests that SMoRA enhances the sharing of related knowledge. Overall, these results demonstrate SMoRA’s strength in fine-grained knowledge sharing and parameter decoupling.